determine if the following are probability distribution (if no, state why)



X=1,2,3,4 and 5



P(x)= 3/10, 1/10, 1/10, 2/10 and 3/10

1
Expert's answer
2022-01-31T15:42:19-0500

By using the conditions of this problem, we may construct a probability space by the following:

1) Take "\Omega=\{1,2,3,4,5\}" as a sample space;

2) Take "\mathbb{F}_{\Omega}=2^{\Omega}", which is the sigma-algebra of all subsets of "\Omega", as an event space;

3) Take "Pr:\mathbb{F}_{\Omega}\to[0,+\infty)", "A\mapsto \sum\limits_{x\in A}P(x)" as a probability function.

This function is additive: for all disjoint subsets "A,B\subset\Omega", we have

"Pr(A\cup B)=\sum\limits_{x\in A\cup B}P(x)=\sum\limits_{x\in A}P(x)+\sum\limits_{x\in B}P(x)"

"=Pr(A)+Pr(B)".

Since "\Omega" is a finite set, the finite additiveness of "Pr" is equivalent to its sigma-additiveness.

Since for all "x\in\Omega", "P(x)>0", the function "Pr" is non-negative.

Since "Pr(\Omega)=\sum\limits_{x\in \Omega}P(x)=P(1)+P(2)+P(3)+P(4)+P(5)"

"=\frac{3}{10}+\frac{1}{10}+\frac{1}{10}+\frac{2}{10}+\frac{3}{10}=1", the total probability is equal to 1.

Therefore, such defined a function "Pr" is indeed a probability function, that is, the function "P(x)" correctly defines a probability distribution on the set {1, 2, 3, 4, 5}.